OpenStudy (ksaimouli):

(fxg)(3) subsitute f(x)=2x-3;g(x)=x+1

OpenStudy (ksaimouli):

plz wxplain

OpenStudy (anonymous):

(fg)(x)=f(x)*g(x)=(2x-3)(x+1)

OpenStudy (anonymous):

(fg)(3)=(2*3-3)((3+1)=3*4=12